I am having trouble completing this word problems for my assignment. I have to use two out of three of the following methods. Substitution, graphic, elimination. Here is the word problem. Quick Copy recently charged $0.49 per page for color pages and $0.07 per page for black-and-white copies. If Shirlee's bill for 90 copies was $11.34, how many copies of each type was made?
Answer by jorel1380(3719)

n+p=90
49n+7p=1134
7n+7p=630
42n=504
n=12
p=78
Shirlee made 12 color copies and 78 black-and-white copies..